OpenStudy (anonymous):

Quadrilateral MATH is a square. If MA = 2x – 5 and AT = x + 10, find the perimeter of the square

OpenStudy (anonymous):

I don't need an answer, I just want to be explained to about how this is done.

OpenStudy (anonymous):

all the sides are equal so 2x-5= x+10

OpenStudy (anonymous):

find x the peri = 4(length of any side)

OpenStudy (smored):

The perimeter of a square is 4 times any side (since every side is congruent to one another), but first you need the length of a side, so 2x-5= x+10. After you solve for x, then x=15. s=x+10, so side=15+10=25. 25*4=100

OpenStudy (anonymous):

;-; that's so simple its not even funny.

OpenStudy (smored):

what do you mean?

OpenStudy (anonymous):

Is this octagon a concave, or a complex?

OpenStudy (anonymous):

Sorry. thank you, I understood how to do it, I needed help on this.

OpenStudy (smored):

concave means that there's an internal angle over 180 degrees (remember conCAVE, there's a hole). a complex polygon has 2 or more sides intersecting itself. Concave Complex